could some help me please
graph each equation
y=3x+2

Answer by jmg(22)

There are two different ways to graph this.
The easiest if you have been working with slope and intercepts is to graph the y-intercept which is the 2. (that means to go up 2 on the graph and put a dot).
Then use the slope, which is 3.
If you write that as a fraction 3/1 you can use rise/run.
This means that you start at the 2 you just graphed and rise or go up 3, and then go to the right or run 1.
So you should have a point graphed at (0,2) and (1, 5). Now just connect the dots.
IF YOU HAVE NOT WORKED WITH SLOPE AND Y-INTERCEPT:
Then make a table of solutions. Another words pick some numbers for x, plug them in and see what you get for y. Then graph the points
Let x=0, then y = 3*0 + 2, so y=2. Graph the ordered pair (0, 2)
Let x = 1, then y = 3*1 + 2 , so y = 5. Graph the ordered pair (1, 5)
Let x = 2, then y = 3*2 + 2, so y = 8. graph the ordered pair (2, 8)
Now try some negatives to get a good representation of the whole graph.
Let x = -1, then y = 3*-1 + 2, so y = -1. Graph (-1, -1)
Let x = -2, then y = 3*-2 + 2, so y = -4. Graph (-2, -4).
Connect the dots and this should give you a pretty accurate representation of the graph.